# Dynamic Adversarial Attacks on Autonomous Driving Systems

### v1

@@ Abstract
This paper introduces an attacking mechanism to challenge the resilience of autonomous driving systems. Specifically, we manipulate the decision-making processes of an autonomous vehicle by dynamically displaying adversarial patches on a screen mounted on another moving vehicle. These patches are optimized to deceive the object detection models into misclassifying targeted objects, e.g., traffic signs. Such manipulation has significant implications for critical multi-vehicle interactions such as intersection crossing and lane changing, which are vital for safe and efficient autonomous driving systems. Particularly, we make four major contributions. First, we introduce a novel adversarial attack approach where the patch is not co-located with its target, enabling more versatile and stealthy attacks. Moreover, our method utilizes dynamic patches displayed on a screen, allowing for adaptive changes and movement, enhancing the flexibility and performance of the attack. To do so, we design a Screen Image Transformation Network (SIT-Net), which simulates environmental effects on the displayed images, narrowing the gap between simulated and real-world scenarios. Further, we integrate a positional loss term into the adversarial training process to increase the success rate of the dynamic attack. Finally, we shift the focus from merely attacking perceptual systems to influencing the decision-making algorithms of self-driving systems. Our experiments demonstrate the first successful implementation of such dynamic adversarial attacks in real-world autonomous driving scenarios, paving the way for advancements in the field of robust and secure autonomous driving.
1 Introduction
@@
Figure 1: Sequential frames demonstrating a successful attack: The top row shows three consecutive frames with the patch displayed on the screen of the patch car. The patch is designed to mislead the camera car’s perception system, i.e., the camera car thinks the Go sign is a Stop sign.
@@
  • We conduct extensive experiments to show successful attacks across various multi-vehicle interaction scenarios including intersection passing and lane changing, and attacking various traffic signs including Go-straight, Turn, and Pedestrian signs. This allows the attacker to adversarially manipulate other vehicles’ decisions to take the lead and move with lower risk.
@@
2 Related Work
In the field of autonomous vehicles, the authors in [43] proposed systematic solutions for Hiding Attack (HA) and Appearing Attack (AA), enhancing the robustness of adversarial examples (AEs) against varying distances and angles. In [14], a method for generating physical AEs that fools traffic sign recognition of autonomous vehicles was presented which employs single and multiple bounding box filters, as well as considering different attack vectors, such as hiding, appearance, non-target, and target attacks. In [18], the authors proposed Short-Lived Adversarial Perturbations (SLAP) that allow adversaries to create physically robust real-world AEs using a projector that provides better control over the attack compared to traditional adversarial patches, as projections can be turned on and off as needed and leave no obvious trace of an attack. Furthermore, in [12], a dynamic adversarial patch technique was introduced to circumvent the object detection model by employing several dynamic patches applied to a target object that adapt to the location of the camera. The authors showed by switching between optimized patches based on the camera’s position, the attack can adapt to different situations and achieve optimal results.
In all the previous physical targeted attacks, the patch is deployed at the location of the target object, which makes the patch implementation hard and even unfeasible in most practical cases. In this paper, we focus on the AEs in which a patch is in a different place from the target object and is dynamically displayed through the screen mounted on a either moving or stationary vehicle. Moreover, the final goal is to mislead the autonomous driving system’s decision-making instead of purely deceiving its perceptions.
3 Problem Formulation
@@ 3.1 Attacker
The variable $\phi$ represents a specific color transformation of the displayed patch in the camera’s image over all possible ones. We discuss it in the Section 4.2.
3.2 Objectives
@@
$\underset{\delta}{\text{maximize}}$ $P[f(\pi(x,\delta,s_{c},s_{p}),\theta)=({\bm{p}}_{\text{target}},{\bm{b}}_{\text{target}})],$ (2)
@@
$\pi(x,\delta,s_{c},s_{p})$ denotes the input image after applying the transformed patch. The patch must be applied within the coordination of the display screen which depends on the positions of the cars. $T(\delta,s_{c},s_{p},\phi)$ denotes the adversarial patch displayed on the screen transformation function provided earlier in Equation 1. ${\bm{p}}_{\text{target}}$ denotes the output prediction corresponding to the confidence of target. ${\bm{b}}_{\text{target}}$ denotes the output prediction corresponding to the bounding box of the target.
We utilize gradient descent optimization to identify the patch that maximizes the misclassification rate of the detector model. Gradient descent is widely used for training adversarial attacks on DNNs [16]. In our case, we aim to optimize the patch that minimizes the loss function, as described in the section 3.3.
According to the proposed function 1 for the patch, it depends on the position of the cars ($s_{p},s_{c}$) and the environmental parameters. In Section 4.1, we introduce how to capture the changes in the positions of the cars and integrate these, resulting in a dynamic patch. To integrate the environmental parameters $\Phi$, we propose a solution in Section 4.2.
3.3 Objective Function
@@
$\delta^{*}=\min_{\delta}-\mathcal{O}(\delta).$ (4)
${\bm{p}}_{\text{target}}(i)$ is the probability of the $i$-th bounding box being the target class (e.g., stop sign). ${\bm{p}}_{\text{obj}}(i)$ is the objectiveness probability of the $i$-th bounding box. The $\text{IoU}({\bm{b}}_{\text{target}}(i),{\bm{b}}_{\text{orig}})$ is the Intersection over Union (IoU) between the $i$-th bounding box of the target class (e.g., stop sign). The ${\bm{b}}_{\text{target}}(i)$ predicted by model and the original object bounding box ${\bm{b}}_{\text{orig}}$. $\mathbb{M}$} is the set of all the bounding boxes overlapping the original object. $\delta^{*}$ is the optimized patch. The objective function, $\mathcal{O}$, is designed to optimize two key factors.
@@
4 Approach
This section elaborates on our proposed process for generating dynamic adversarial patches. Section 4.1 explains how the patch training data is collected capturing the dynamic of the setting. In Section 4.2, we encounter the image transformation of the displayed patch on the screen. Finally, we describe the patch optimization in Section 4.3.
4.1 Dynamic Environment Settings
The predictions of object detectors are influenced by the positions of objects in the image. To visualize the principal components of the learned features and representations, we apply Eigen-CAM [23, 11] which reflects the influence level of each image region on the model through a heatmap (red to purple denoting most to least influence). As depicted in Figure 2, the model’s influence shifts according to changes in object positions. This in turn impacts the effectiveness of adversarial patches that are not overlaid on the target object. To compensate, a dynamic patch that adapts to the target’s position could be a more effective approach. Without direct access to the camera car’s captured images during an attack, the patch car can estimate the target’s position and adjust its patch accordingly to account for positional changes.
@@ 4.1 Dynamic Environment Settings
4.1.1 Data collection
@@
4.1.2 Data clustering
The recorded data is then clustered by the K-means clustering algorithm based on the distance of the patch car and target object from the camera car. This way, the positions of the patch car and the target in the images in the same cluster are within the same threshold Each cluster represents a specific range of distances, and a dedicated patch is trained for each cluster. Figure 3(b) shows the clustered points of the patch car and the target. This enhances the performance of the patches that are optimized for different.
@@ 4.1.2 Data clustering
This structured and strategic approach to data collection and processing ensures that our system is equipped to handle a variety of real-world scenarios, making the patches dynamic and enhancing the overall effectiveness of the visual transformations.
(b) Loss over training epochs for SIT-Net.
Figure 4: SIT-Net predictions and the loss during its training.
4.2 Screen Image Transformation Network (SITNet)
@@
4.2.1 Screen Data Collection and Prepossessing
@@
Input image generation. A diverse set of patch-like images was created, ensuring a variety of color and contrast scenarios. These patches are of the same size as the targeted screen area to maintain consistency.
Screen display. The generated images were displayed on a screen, capturing a wide range of total variations to adequately represent different potential real-world conditions.
Image recording. A camera was utilized to record the displayed images. Precise coordination of the four corner points of the screen area was maintained to aid in the subsequent image-processing steps.
Image preprocessing. The captured images were subjected to cropping and perspective transformation based on the known screen coordinates. This step ensures that the images are accurately aligned and standardized, mirroring the conditions of the screen display.
@@
4.2.2 Loss Function and Optimization
The mean squared error is used in the loss function to reduce the difference between predicted and real-world images. However, these pixel-level differences might not correlate well with how images are perceived or their similarity. Perceptual loss addresses this limitation by incorporating features extracted from pre-trained neural networks that are designed to mimic human visual perception. Moreover, a total variation term [21] is added to the loss function to smooth the predicted images.
Mean squared error (MSE) loss. We quantify the pixel-wise difference between the predicted and real-world images by MSE loss.
@@ 4.2.2 Loss Function and Optimization
Perceptual loss. The perceptual loss measures the difference in style between images at the feature level rather than the pixel level. The first nice layers of the pre-trained VGG19 network [29] are used to extract feature maps from the input and target images. Then the MSE between these feature maps is computed. This way, the model is forced to generate images that are perceptually similar to the target images. The perceptual loss is defined below.
@@
Figure 5: The patch optimization pipeline.
4.3 Patch Optimization
The optimization procedure for the patch is described in Algorithm 1. We optimize a patch for each cluster in the dataset. First, the patch is transformed by the SIT-Net. Then we apply the transformed patch, $z^{\prime}$, to the images and pass it to the object detector. The line 7 includes filtration of the object detector’s prediction and extraction of all bounding boxes in the area of the original object. The objectiveness probabilities of these bounding boxes are integrated into the class probabilities. Then we extract the top-$k$ target class confidence values.
@@ 4.3 Patch Optimization
0: Set of clustered images ${\mathbb{X}}_{\text{clustered}}$, original traffic sign bounding box ${\bm{b}}_{\text{orig}}$, patch $z\in{\mathbb{P}}^{\prime}$, learning rate $\eta$, number of iterations NN, top-$k$ value $k$, SITNet model $M_{\text{SITNet}}$
@@
Subsequently, in line 9, the IoU of the top-$k$ target class bounding boxes with the original object bounding box is computed. In the next step, the IoU values are multiplied by the top-$k$ confidence, and its mean is calculated. This value represents the alignment and confidence of the selected bounding boxes. Finally, This value is utilized in the optimization process of the patch. Figure 5 shows a pipeline of the patch optimization process.
5 Evaluation
@@
(a) The intersection scenario
(b) The lane changing scenario
Figure 6: Experimental setup. Left column: images recorded by the camera car; Middle column: point clouds representing the path car and the target sign within the map; Right column: the positioning of the cars.
5.1 Experimental Setup
Agents. We use two Rosmaster R2 robot cars with Ackermann structure [41] equipped with Lidar and IMU sensors for localization and mapping. The camera car uses a Zed2 camera [31] for image recording. The patch car is equipped with a seven-inch LCD screen attached to the side of the robot to display the patches. The car’s decision-making planner employs object detection results to plan for their motions. Object detection and patch training are carried out by a server using a Navidia RTX4090 GPU.
Object detector model. For each robot car, we fine-tuned the model from yolov5 [15] to detect traffic signs and the other robot car. The model trained over 1.8 K images with the weights from a pre-trained model yolov5s and the input image size of $640\times 640$. The dataset is available online. [^1] https://universe.roboflow.com/r2-5io2k/rosmaster-r2-traffic-sign
(a) Loss for the ‘Go-straight’ sign attack
(b) Loss for the ‘Turn’ sign attack
(c) Loss for the ‘Pedestrian’ sign attack
@@ 5.1 Experimental Setup
Autonomous driving scenarios. We consider two different robot positional settings at the intersections where the traffic signs are crucial for decision-making. When the patch car must yield to the camera car according to the traffic signs, it aims to alter the class of the detected signs by the camera car to pass the intersection or change its lane earlier. The first setting is a passing intersection where both cars approach an intersection within the same time window while driving on crossed roads, and the patch car tries to pass the intersection before the camera car. There are three possible traffic signs—Go-straight, Turn, and Pedestrian crossing signs, that are targeted by the attack to be misclassified as a stop sign. In the second setting, both cars are driving on the same road and the patch car tries to merge to the lane of the camera car with lower risk by deceiving the camera car to detect the sign as a stop sign. Figure 6 shows our setup for the experiment.
Near Mid Far All
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Visual representation of attack across different clusters and the entire dataset. Rows correspond to ‘Go-straight’, ‘Turn’, and ‘Pedestrian’ sign scenarios, respectively. Columns distinguish between individual clusters (‘Near’, ‘Mid’, and ‘Far’) and patches trained on the entire dataset (All). This illustrates the adaptive capability of the dynamic patches to manipulate the camera car’s object detection under varying environmental contexts.
Figure 9: Display of trained adversarial patches across various clusters and the entire dataset. The rows represent different traffic sign scenarios, specifically ‘Go-straight’, ‘Turn’, and ‘Pedestrian’ signs, in that order. For the columns, from left to right, we showcase patches optimized for clusters based on proximity to the target: ‘Near’, ‘Mid’, and ‘Far’, followed by patches trained on the combined dataset (denoted as ‘All’). This arrangement illustrates how the patches are tailored to specific distances from the target object.
Method Go-straight sign Turn sign Pedestrian sign
Similar Env. Unseen Env. Similar Env. Unseen Env. Similar Env. Unseen Env.
Dynamic Patch (ours) $\approx$ 28.8% 4.0% 47.5% 5.1% 41.0% 5.0%
Static Patch (SoA) $\approx$ 22.0% 2.0% 32.8% 2.0% 40.8% 7.0%
Total frames 59 98 61 98 61 99
Table 1: Attack success rate of our Dynamic Patch and the state-of-the-art Static Patch for each sign in the Intersection.
5.2 Attack Procedure

### v2

@@ Abstract
This paper introduces an attacking mechanism to challenge the resilience of autonomous driving systems. Specifically, we manipulate the decision-making processes of an autonomous vehicle by dynamically displaying adversarial patches on a screen mounted on another moving vehicle. These patches are optimized to deceive the object detection models into misclassifying targeted objects, e.g., traffic signs. Such manipulation has significant implications for critical multi-vehicle interactions such as intersection crossing, which are vital for safe and efficient autonomous driving systems. Particularly, we make four major contributions. First, we introduce a novel adversarial attack approach where the patch is not co-located with its target, enabling more versatile and stealthy attacks. Moreover, our method utilizes dynamic patches displayed on a screen, allowing for adaptive changes and movements, enhancing the flexibility and performance of the attack. To do so, we design a Screen Image Transformation Network (SIT-Net), which simulates environmental effects on the displayed images, narrowing the gap between simulated and real-world scenarios. Further, we integrate a positional loss term into the adversarial training process to increase the success rate of the dynamic attack. Finally, we shift the focus from merely attacking perceptual systems to influencing the decision-making algorithms of self-driving systems. Our experiments demonstrate the first successful implementation of such dynamic adversarial attacks in real-world autonomous driving scenarios, paving the way for advancements in the field of robust and secure autonomous driving.
I Introduction
@@
(a) Adversarial patch attacks on the Pedestrian sign.
(b) Benign trial.
Figure 1: Sequential frames of a multi-robot interaction in an intersection: The top row of (a) demonstrates a dynamic adversarial patch attack, where the patch is displayed on the screen of the patch car. The patch is designed to mislead the camera car’s perception system, i.e., the camera car thinks the Pedestrian sign is a Stop sign. The top row of (b) illustrates the system’s correct behavior during a benign trial without adversarial interference. The bottom rows of (a) and (b) are the top views of the settings.
@@
  • We conduct extensive experiments to show successful attacks across various multi-vehicle interaction scenarios passing the intersection, and attacking various traffic signs including Go-straight, Turn, and Pedestrian signs. This allows the attacker to adversarially manipulate other vehicles’ decisions to take the lead and move with lower risk.
@@
II Related Work
In the field of autonomous vehicles, the authors in [43] proposed systematic solutions for Hiding Attack (HA) and Appearing Attack (AA), enhancing the robustness of adversarial examples (AEs) against varying distances and angles. In [14], a method for generating physical AEs that fools traffic sign recognition of autonomous vehicles was presented which employs single and multiple bounding box filters, as well as considering different attack vectors, such as hiding, appearance, non-target, and target attacks. In [18], the authors proposed Short-Lived Adversarial Perturbations (SLAP) that allow adversaries to create physically robust real-world AEs using a projector that provides better control over the attack compared to traditional adversarial patches, as projections can be turned on and off as needed and leave no obvious trace of an attack. Furthermore, in [12], a dynamic adversarial patch technique was introduced to circumvent the object detection model by employing several dynamic patches applied to a target object that adapts to the location of the camera. The authors showed by switching between optimized patches based on the camera’s position, the attack can adapt to different situations and achieve optimal results.
In all the previous physical targeted attacks, the patch is deployed at the location of the target object, which makes the patch implementation hard and or impossible in most practical cases. In this paper, we focus on the AEs in which a patch is in a different place from the target object and is dynamically displayed through the screen mounted on a either moving or stationary vehicle. Moreover, the final goal is to mislead the autonomous driving system’s decision-making instead of purely deceiving its perceptions.
III Problem Formulation
@@ III-A Attacker
The variable $\phi$ represents a specific color transformation of the displayed patch in the camera’s image over all possible ones. We discuss it in the Section IV-B.
III-B Objectives
@@
[…]
@@
$\pi(x,\delta,s_{c},s_{p})$ denotes the input image after applying the transformed patch. The patch must be applied within the coordination of the display screen which depends on the positions of the cars. $T(\delta,s_{c},s_{p},\phi)$ denotes the adversarial patch displayed on the screen transformation function provided earlier in Equation 1. ${\bm{p}}_{\text{target}}$ denotes the output prediction corresponding to the confidence of the target. ${\bm{b}}_{\text{target}}$ denotes the output prediction corresponding to the bounding box of the target.
We utilize gradient descent optimization to identify the patch that maximizes the misclassification rate of the detector model. Gradient descent is widely used for training adversarial attacks on DNNs [16]. In our case, we aim to optimize the patch that minimizes the loss function, as described in the section III-C.
According to the proposed function 1 for the patch, it depends on the position of the cars ($s_{p},s_{c}$) and the environmental parameters. In Section IV-A, we introduce how to capture the changes in the positions of the cars and integrate these, resulting in a dynamic patch. To integrate the environmental parameters $\Phi$, we propose a solution in Section IV-B.
III-C Objective Function
@@
$\delta^{*}=\min_{\delta}-\mathcal{O}(\delta).$
${\bm{p}}_{\text{target}}(i)$ is the probability of the $i$-th bounding box being the target class (e.g., stop sign). ${\bm{p}}_{\text{obj}}(i)$ is the objectiveness probability of the $i$-th bounding box which is a likelihood that measures the confidence that the bounding box contains an object. The $\text{IoU}({\bm{b}}_{\text{target}}(i),{\bm{b}}_{\text{orig}})$ is the Intersection over Union (IoU) between the $i$-th bounding box of the target class (e.g., stop sign). The ${\bm{b}}_{\text{target}}(i)$ is predicted by model and ${\bm{b}}_{\text{orig}}$ is the original object bounding box. ${\mathbb{M}}$ is the set of all the bounding boxes overlapping the original object. $\delta^{*}$ is the optimized patch. The objective function, $\mathcal{O}$, is designed to optimize two key factors:
@@
IV Approach
This section elaborates on our proposed approach for generating dynamic adversarial patches. Section IV-A explains the method for collecting patch training data that captures the dynamics of the environment. In Section IV-B, we demonstrate the image transformation of the patch that is displayed on the screen. Finally, we describe the patch optimization in Section IV-C.
IV-A Dynamic Environment Settings
The output of the object detection models (e.g., the regular yolov5s [15]) such as cars or traffic signs can be greatly influenced by the objects’ position in the image. This positioning directly impacts the model’s focus or “attention”, altering the detection outcomes based on how far or near these objects are positioned relative to the camera’s perspective. To visualize the principal components of the learned features and representations, we apply Eigen-CAM [23, 11] which reflects the influence level of areas of an image on the model through a heatmap (red to purple denoting most to least influence). As depicted in Figure 2, the model’s attention shifts according to changes in object positions. This in turn impacts the effectiveness of adversarial patches that are not overlaid on the target object. To compensate, a dynamic patch that adapts to the target’s position could be a more effective approach. Without direct access to the camera car’s captured images during an attack, the patch car can estimate the target’s position and adjust its patch accordingly to account for positional changes.
(b) Loss over training epochs for SIT-Net.
Figure 4: SIT-Net predictions and the loss during its training.
IV-A1 Data collection
@@
IV-A2 Data clustering
The recorded data is then clustered by the K-means clustering algorithm based on the distance of the patch car and target object from the camera car. This way, the positions of the patch car and the target in the images in the same cluster are within the same threshold. Each cluster represents a specific range of distances, and a dedicated patch is trained for each cluster. Figure 3(b) shows the clustered points of the patch car and the target. This enhances the performance of the patches that are optimized for different clusters. Appendix A-A provides a more detailed overview of our dataset.
@@ IV-A2 Data clustering
This structured and strategic approach to data collection and processing ensures that our system is capable of handling a variety of real-world scenarios, making the patches dynamic and enhancing the overall effectiveness of the visual transformations.
Figure 5: The patch optimization pipeline.
IV-B Screen Image Transformation Network (SITNet)
@@
IV-B1 Screen Data Collection and Prepossessing
@@
Input image generation. A diverse set of patch-like images is created, ensuring a variety of color and contrast scenarios. These patches are of the same size as the targeted screen area to maintain consistency.
Screen display. The generated images are displayed on a screen, capturing a wide range of total variations to adequately represent different potential real-world conditions.
Image recording. A camera is utilized to record the displayed images. Precise coordination of the four corner points of the screen area is maintained to aid in the subsequent image-processing steps.
Image preprocessing. The captured images are subjected to cropping and perspective transformation based on the known screen coordinates. This step ensures that the images are accurately aligned and standardized, mirroring the conditions of the screen display.
@@
IV-B2 Loss Function and Optimization
The mean squared error is used in the loss function to reduce the difference between predicted and real-world images. However, these pixel-level differences might not correlate well with how images are perceived or their similarity. Perceptual loss addresses this limitation by incorporating features extracted from pre-trained neural networks that are designed to mimic human visual perception. Moreover, a total variation term [21] is added to the loss function to smooth the predicted images. Mean squared error (MSE) loss. We quantify the pixel-wise difference between the predicted and real-world images by MSE loss.
@@ IV-B2 Loss Function and Optimization
Perceptual loss. The perceptual loss measures the difference in style between images at the feature level rather than the pixel level. The first nine layers of the pre-trained VGG19 network [29] are used to extract feature maps from the input and target images. Then the MSE between these feature maps is computed. This way, the model is forced to generate images that are perceptually similar to the target images. The perceptual loss is defined below.
@@
0:  Set of clustered images ${\mathbb{X}}_{\text{clustered}}$, original traffic sign bounding box ${\bm{b}}_{\text{orig}}$, patch $z\in{\mathbb{P}}^{\prime}$, learning rate $\eta$, number of iterations $N$, top-$k$ value $k$, SITNet model $M_{\text{SITNet}}$
@@
IV-C Patch Optimization
The optimization procedure for the patch is described in Algorithm 1. We optimize a patch for each cluster in the dataset. First, the patch is transformed by the SIT-Net. Then we apply the transformed patch, $z^{\prime}$, to the images and pass it to the object detector. The line 7 includes filtering of the object detector’s prediction and extraction of all bounding boxes in the area of the original object. The objectiveness probabilities of these bounding boxes are integrated into the class probabilities. Then we extract the top-$k$ target class confidence values. Subsequently, in line 9, the IoU of the top-$k$ target class bounding boxes with the original object bounding box is computed. In the next step, the IoU values are multiplied by the top-$k$ confidence, and its mean is calculated. This value represents the alignment and confidence of the selected bounding boxes. Finally, this value is utilized in the optimization process of the patch. Figure 5 shows a pipeline of the patch optimization process. In summary, this process begins with the clustering of the images through the K-Means method using LiDAR data, then the patches that are transformed by the SIT-Net are applied to images of each cluster. The patches are optimized by backpropagating.
V Evaluation
@@
V-A Experimental Setup
Agents. We use two Rosmaster R2 robot cars with Ackermann structure [41] equipped with Lidar and IMU sensors for localization and mapping. The camera car uses a Zed2 camera [31] for image recording. The patch car is equipped with a seven-inch LCD screen attached to its side to display the patches. The car’s decision-making planner employs object detection results to plan for its motions. Object detection and patch training are carried out by a server using a NVIDIA RTX4090 GPU.
Object detector model. For each robot car, we fine-tuned the model from yolov5 [15] to detect traffic signs and the other robot car. The model trained over 1.8 K images with the weights from a pre-trained model yolov5s and the input image size of $640\times 640$. Our customized dataset is attached. [^1] The link to dataset is provided within supplementary materials.
Figure 6: Experimental setup. Left column: images recorded by the camera car; Middle column: point clouds representing the patch car and the target sign within the map; Right column: the top view of the two cars.
(a) Loss for the “Go-straight” sign attack
(b) Loss for the “Turn” sign attack
(c) Loss for the “Pedestrian” sign attack
@@ V-A Experimental Setup
Autonomous driving scenarios. We consider various robot positional settings at the intersections where the traffic signs are crucial for decision-making. When the patch car must yield to the camera car according to the traffic signs, it aims to alter the class of the detected signs by the camera car to pass the intersection earlier. In this setting, both cars approach an intersection within the same time window while driving on crossed roads, and the patch car tries to pass the intersection before the camera car. There are three possible traffic signs: Go-straight, Turn, and Pedestrian crossing signs. The sign is targeted by the patch attack to be misclassified as a stop sign. Figure 6 shows our setup for the experiment.
V-B Attack Procedure